OpenStudy (anonymous):

Dylan opened a credit card account with $650.00 of available credit. Now that he has made some purchases, Dylan's account only has $546.00 of available credit. What is the percentage decrease of the amount of available credit in Dylan's account?

OpenStudy (anonymous):

OhKay, do you know how to find the percentage?

OpenStudy (anonymous):

Yes, I know how to find the percentage.

OpenStudy (anonymous):

Then, alright. In order to find the decrease of the percentage you need to subtract 650-546.

OpenStudy (anonymous):

650-546=????

OpenStudy (anonymous):

Sorry about that. 650-546=104. Correct?

OpenStudy (anonymous):

Yes, now we know that there is 104$ less. now you divide 104/605=??? than after that you multiply it by 100 so, 100*???= (your answer)

OpenStudy (anonymous):

Sorry, 104/650

OpenStudy (anonymous):

Ohhh. 104/650=0.16 0.16*100=16...?

OpenStudy (anonymous):

It 16% decrease.c:

OpenStudy (anonymous):

Oh. Haha. I forgot to add the percentage sign xD Thank you so so so so much!